exact Find the value of cos(x + y). Given sin x = žand cos y = 5, x is an angle in Quadrant II and y is an angle in QuadrantMay you show some steps? I am completely clueless on how to solve these type of questions. Thank you!
Answer #1

A that given sina = 3 1 = 3/5 5 3 X by using pythagorean theorem %= a +6 s =ăt 3 5²_3 25-9 16 = 4 adjacent SDAY a= an. a= Il